# Standing on the Shoulders of Giants: Reprogramming Visual-Language Model for General Deepfake Detection

### v1

@@ Abstract
The proliferation of deepfake faces poses huge potential negative impacts on our daily lives. Despite substantial advancements in deepfake detection over these years, the generalizability of existing methods against forgeries from unseen datasets or created by emerging generative models remains constrained. In this paper, inspired by the zero-shot advantages of Vision-Language Models (VLMs), we propose a novel approach that repurposes a well-trained VLM for general deepfake detection. Motivated by the model reprogramming paradigm that manipulates the model prediction via data perturbations, our method can reprogram a pre-trained VLM model (e.g., CLIP) solely based on manipulating its input without tuning the inner parameters. Furthermore, we insert a ”pseudo-word” guided by facial identity into the text prompt. Extensive experiments on several popular benchmarks demonstrate that (1) the cross-dataset and cross-manipulation performances of deepfake detection can be significantly and consistently improved (e.g., over 88% AUC in cross-dataset setting from FF++ to WildDeepfake) using a pre-trained CLIP model with our proposed reprogramming method; (2) our superior performances are at less cost of trainable parameters, making it a promising approach for real-world applications.
@@ Introduction
Deepfake refers to a series of deep learning-based facial forgery techniques(Li et al. 2020a; Yao et al. 2021; Xu et al. 2022; Chen et al. 2024) that can swap or reenact the face of one person in a video to another. In recent years, deepfake videos (a.k.a, deepfakes) have gained substantial attention due to their potential by creating and spreading false information. Thus, detecting deepfakes has emerged as a crucial research topic to reduce such security risks.
Existing methods treat deepfake detection as a binary classification problem and predominantly utilize CNNs (e.g., Xception (Chollet 2017) or EfficientNet (Tan and Le 2019)) as the backbones of classifier. In addition, some works(Qian et al. 2020; Li et al. 2020b; Zhao et al. 2021; Liu et al. 2021) propose to introduce auxiliary clues, including modalities (e.g., frequency) or supervision (e.g., forgery masks) information for learning subtle forgery artifacts. Despite these advancements, evaluating forgeries from unseen datasets and synthesized by unseen methods beyond the training data still poses a significant challenge for practical deepfake detection.
@@
To enhance the generalizability of deepfake detectors, several works employ vision transformer (ViT)(Dosovitskiy et al. 2021) models as the backbone, owing to their superior image representation capabilities compared to CNNs. Nevertheless, ViT-based detection methods, e.g. UiA-ViT(Zhuang et al. 2022) and Bi-LIG(Jiang et al. 2024), typically involve fully fine-tuning pre-trained ViT-based backbones (e.g., with ImageNet weights). This process demands extensive computational and storage resources, which hinders their practical deployment in real-world applications. Thus, developing a general yet cost-efficient approach is necessary for deepfake detection.
Due to the inherent vulnerability of deep networks to adversarial attacks, a novel concept called model reprogramming (Elsayed, Goodfellow, and Sohl-Dickstein 2019; Chen 2024) has been proposed to re-purpose a well-trained model trained in a source domain to perform a target-domain task through learning a universal perturbation without modifying the source-domain model parameters. In other words, an additive offset applied to a deep network’s input would be sufficient to adapt the network to a new task without the need of re-training or fine-tuning its inner parameters. It is evident that the reprogramming paradigm requires significantly fewer computing resources compared to fine-tuning numerous parameters within the model. Inspired by the above advancements, we advocate that reprogramming a well-trained foundation model to identify deepfakes is a promising way in terms of computational efficiency. Considering the Vision Language Models (VLMs) (Bordes et al. 2024) have shown impressive zero-shot performances by bridging vision and language modalities, we employ CLIP (Radford et al. 2021) as the foundation model and adapt it to identify deepfakes.
To this end, in this paper, we propose RepDFD, a novel method to reprogram a pre-trained CLIP model for effective and general deepfake detection. RepDFD solely learns single perturbations in pixel spaces for the deepfake detection task while keeping the entire CLIP model frozen. Specifically, we introduce Input Transformation to merge the image and learnable perturbations and then feed it into the image encoder of CLIP. Furthermore, we propose Face2Text Prompts to generate text prompts merging facial identity information, and then feed them into the text encoder of CLIP to guide the optimization of perturbations. During the testing phase, the sole operation involves adding such task-specific perturbations (a.k.a, visual prompts) to all test images before feeding them into the CLIP model for evaluation. This straightforward method enables the CLIP model to effectively detect deepfakes. Moreover, since the internal parameters are not trained, the foundation CLIP model can be reused for the other vision tasks. Extensive experiments on several popular deepfake benchmarks demonstrate that (1) the cross-dataset and cross-manipulation performances can be significantly and consistently enhanced by equipping RepDFD for a pre-trained CLIP model; and (2) our superior performances are at less cost of trainable parameters (See in Figure 1), making it a promising approach for real-world applications.
@@
  • This is the first work to explore utilizing the model reprogramming paradigm for deepfake detection.
  • We propose novel RepDFD to reprogram a pre-trained CLIP model by solely processing its image and text inputting without tuning the inner parameters. Thus, RepDFD can be seamlessly adapted to other foundation models in a plug-and-play way.
  • We conduct extensive experiments on several benchmarks, and demonstrate that RepDFD is general and efficiency for deepfake detection.
Related Works
@@
The past five years have witnessed a wide variety of methods proposed for defending against the malicious usage of deepfakes. Currently, the majority of deepfake detection methods are based on deep learning, leveraging generic CNNs (e.g., Xception(Chollet 2017), EfficientNet(Tan and Le 2019)) as the backbones of the classifier. Furthermore, several works (Qian et al. 2020; Liu et al. 2021; Zhao et al. 2021) utilize frequency information or localize the forged regions to improve the performance of detectors. Nevertheless, the generalization challenge still hinders the application of deepfake detectors in real-world scenarios. To address such issue, several works (Li et al. 2020b; Shiohara and Yamasaki 2022; Larue et al. 2023; Nguyen et al. 2024) introduce the pseudo deepfake data, which belnding two different faces, during the training. However, all of the above methods typically involve retraining at least one backbone network. In general, employing more powerful backbone networks (e.g., replacing CNNs with ViTs) produces better performances but at the cost of increased computational resources.
@@
CLIP is a vision-language model (Radford et al. 2021), that is able to perform flexible zero-shot transfer to unseen classes using text prompts. Since CLIP is pre-trained to predict whether an image matches a textual description, it naturally fits zero-shot recognition. This is achieved by comparing image features with the classification weights synthesized by the text encoder, which takes as input textual descriptions specifying classes of interest. The CLIP model contains an image-encoder $E_{I}(\cdot)$ and a text-encoder $E_{T}(\cdot)$ such that the cosine similarity between the features $E_{I}(x_{k})$ and $E_{T}(t_{k})$ are maximized with respect to each pair $k$. Compared with the traditional classifier learning approach where closed-set visual concepts are learned from random vectors, vision-language pre-training allows open-set visual concepts to be explored through a high-capacity text encoder, leading to a broader semantic space and in turn making the learned representations more transferable to various vision tasks. Several works(Sha et al. 2023; Ojha, Li, and Lee 2023; Cozzolino et al. 2024) has explore that using CLIP model for detecting generated image by GANs and Diffusion Models. VLFFD proposes to fully fine-tune CLIP with fine-grained text prompts for deepfake detection. In this work, we propose a cost-efficient method to adapt a large pre-trained CLIP model for general deepfake detection.
@@
Figure 2: Overall framework of our proposed method. The core idea involves optimizing an universal visual prompt on a frozen CLIP model to adapt it for the deepfake detection task.
Parameter-Efficient Tuning
Parameter Efficient Fine-Tuning (PEFT) methods, which significantly reduce computational and storage overheads, have garnered increasing attention. By fine-tuning only a limited number of additional parameters, PEFT can adapt a large pretrained model to achieve outstanding performance in targeted tasks.
@@
One of the PEFT methods, Model Reprogramming (Chen 2024), which incorporates prompts into inputs, offers an effective framework for repurposing models for various task-specific applications. The framework draws significant inspiration from adversarial reprogramming, which was first introduced by Elsayed et al (Elsayed, Goodfellow, and Sohl-Dickstein 2019) The aim of model reprogramming is to re-use and re-align the data representation, from an existing model, for a separate task without fundamental changes to the model’s parameters. This paradigm re-purposes existing knowledge by strategically transforming inputs and outputs, bypassing extensive inner model parameter fine-tuning. Reprogramming techniques have been widely applied in various tasks in the past few years. VP introduces a universal perturbation directly into the input data to facilitate task-specific fine-tuning while keeping the pretrained model intact. Although these PEFT methods have developed in other research fields, the application of PEFT methods to face forgery detection remains largely unexplored. In this study, we introduce a novel approach to reprogramming a pre-trained CLIP model for deepfake detection using a minimal number of learnable parameters, designed to leverage the great generalization abilities of the pretrained Vision-Language Model (VLM).
@@
The goal of our method is to adapt the frozen CLIP model for deepfake detection task by solely learning a single visual prompt. Formally, given a well-trained CLIP model (including the image encoder $E_{I}(\cdot)$ and the text encoder $E_{T}(\cdot)$), we learn a single visual prompt $\boldsymbol{\delta}$ for adapting the frozen $E_{I}(\cdot)$ and $E_{T}(\cdot)$ to identify deepfakes. To achieve this, we propose Input Transformation and Face2Text Prompts, which solely process the image and text before inputting $E_{I}(\cdot)$ and $E_{T}(\cdot)$, respectively. During evaluation, the optimized $\boldsymbol{\delta}$ is added to the input images to enhance the visual clues of identifying deepfakes. In this way, we harness the power of learned perturbations to guide the CLIP model to focus the deepfake detection task while preserving its inner parameters. The overall framework is depicted in Figure 2. In what follows, we elaborate on the details of Input Transformation, Face2Text Prompts and the optimization pipeline.
@@
In designing the placement of visual prompts, we adopt a similar scheme to (Bahng et al. 2022; Tsao et al. 2024), wherein visual prompts are placed around images for input transformations. Although the input image size is determined by the source model, when fine-tuning to a downstream dataset from a source model, there is design freedom to resize the target images and use the remaining space for visual prompts. Figure 3 demonstrates the details of Input Transformation. Given an original image $\mathbf{X}$, the input transformation aims to merge the universal perturbation $\boldsymbol{\delta}$ to $\mathbf{X}$. Mathematically, we describe the process of input transformations as
$\mathcal{T}_{p}(\mathbf{X},\boldsymbol{\delta})=\text{Resize}_{p}(\mathbf{X})+\mathbf{M}_{p}\odot\boldsymbol{\delta}$ (1)
where $p$ is the width of the visual prompt and $M_{p}$ is a binary mask with a border of width $p$. The original image $X$ of size $H\times W$ is resized to $H^{\prime}\times W^{\prime}=(H-2p)(W-2p)\rfloor$, so that the prompted image is of the same size as the original without overlapping with $\boldsymbol{\delta}$. Note that the number of trainable parameters in our method is model-agnostic, which can be computed by
#$\#Para=3(HW-H^{\prime}W^{\prime})=6p(H+W)-12p^{2}$ (2)
This expression indicates that $\boldsymbol{\delta}$ is applied to the RGB channels of the image.
We feed $\mathcal{T}_{p}(\mathbf{X},\boldsymbol{\delta})$ into image encoder to get the image feature $f$, i.e.,
$f=E_{I}(\mathcal{T}_{p}(\mathbf{X},\boldsymbol{\delta}))$ (3)
@@
To map model outputs to the target label, the CLIP model utilizes text prompts to align image features in a shared latent space, enabling classification by matching the closest text description to the image features. For instance, one of the official unified templates of text prompts is “A photo of a [cls]”, where [cls] is a placeholder of class label. In the context of deepfake detection, the binary class labels “real face” and “fake face” (or synonymous terms) lack specificity compared to labels such as “cat” or “dog”, making them difficult to accurately identify using existing text templates. To address this issue, VLFFD (Sun et al. 2023) proposes to introduction of fine-grained text descriptions manually to fine-tune both $E_{I}(\cdot)$ and $E_{T}(\cdot)$ in the CLIP model. However, designing a hand-crafted text prompt that comprehensively describes forgery clues is difficult and labor-intensive.
In this work, we aim to design simple yet effective text prompts to fully harness the generalizable capabilities of the CLIP model for deepfake detection. The success of personalized image generation methods has shown that introducing pre-trained image embedding as textual embeddings is beneficial for maintaining the consistency of generated subjects. Considering that deepfakes involve the manipulation of faces, the inconsistency of facial representations between real and fake faces is a reliable clue. Inspired by this, we suggest exploit the CLIP text encoder to integrate face embeddings into text embeddings. Therefore, the face information, which is difficultly expressed by language, can be incorporated into textual prompts, thereby enhancing the CLIP model’s capability for efficient deepfake detection.
To achieve this, a Face2Text module is proposed and the details are shown in Figure 4. The prompt template is designed as “A [cls] photo of a [ID] person”. Given a face encoder $F_{id}$, we represent a face embedding $S^{*}$ for an input image $\mathbf{X}$:
$\centering S^{*}=f_{\text{map}}(F_{\text{id}}(\mathbf{X})).\@add@centering$ (4)
Following tokenization, we substitute the placeholder token [ID] with the face embedding $S^{*}$, yielding a sequence of token. $f_{map}$ represent a random projection layer to align the dimension. Herein, we employ a pre-trained Transface (Dan et al. 2023) as $F_{id}$.
@@
We also consider the plaint text template that “A [cls] photo of a person”. Thus, we get four different text template for real and fake labels (See in Table1) One of interesting finding according to results the is that the class-specific setting $\mathbf{T}=\{T_{0},T_{3}\}$ achieves the best performance. The in-depth analysis are discussion later.
We feed tokenized $\mathbf{T}$ into text encoder to get the text feature $w$, i.e.,
$w=E_{T}(ToEmbedding(\mathbf{T})),$ (5)
where the operation $ToEmbedding$ represents the process of transforming words into word embeddings.
@@
$T_{0}$ “A real photo of person”
$T_{1}$ “A fake photo of person”
$T_{2}$ “A real photo of [ID] person”
$T_{3}$ “A fake photo of [ID] person”
@@
$\centering P(y=i\mid x)=\frac{\exp(\cos(w_{i},f)/\tau)}{\sum_{j=1}^{2}\exp(\cos(w_{j},f)/\tau)},\@add@centering$ (6)
where $cos(\cdot)$ denotes the cosine similarity and $\tau$ is the temperature parameter of CLIP. Eq. (6) have consistent formulation to a traditional full connected layer with input dimension $N$ and output dimension $2$ if we treat the $[w_{r},w_{f}]$ as the ‘weight’ of a full connected layer.
During training we optimize the prompt by maximizing the likelihood of the correct label, and during inference we pad the optimized prompt around the shrunk test samples for predictions. For a training dataset $\mathcal{D}$, the optimization target can be formulated as minimizing $\mathcal{L}$ as
$\centering\mathcal{L}=\mathbb{E}_{\mathbf{X}\sim\mathcal{D}}\left[-\sum_{i}^{N}P(y_{i}|x)\log(P(y_{i}|x))\right],\@add@centering$ (7)
We initialize the $\boldsymbol{\delta}$ to zero, and follow a simple gradient-based approach to directly optimize the visual prompt via back-propagation which updates $\boldsymbol{\delta}$ at step $t$ by
$\centering\boldsymbol{\delta}^{t+1}=\boldsymbol{\delta}^{t}-\gamma{\nabla_{\boldsymbol{\delta}^{t}}\mathcal{L}},\@add@centering$ (8)
@@
SLADD (CVPR 2022) 79.70 - - 77.20 SBI (CVPR 2022) 93.18 - 86.15 72.42
@@
Following most previous works, we mainly conducted experiments on the FaceForensics++ (FF++)(Rossler et al. 2019). It con-tains 1000 Pristine (PT) videos (i.e., the real sample) and 5000 fake videos forged by five manipulation methods, i.e., Deepfakes (DF), Face2Face (F2F), FaceSwap (FS), NeuralTextures (NT) and FaceShifter (FSh). Besides, FF++ provides three quality levels in compression for these videos: raw, high-quality (HQ) and low-quality (LQ). The HQ version of FF++ is adopted by default in this paper. The samples were split into disjoint training, validation, and testing sets at the video level follows the official protocol.
To demonstrate the performances in cross-dataset settings, four additional datasets are adopted, i.e., Celeb-DF-v2 (CDF) (Li et al. 2020c), DeepFake Detection Challenge preview (DFDCP) (Dolhansky et al. 2019), DeepFake Detection Challenge public (DFDC) (Dolhansky et al. 2020) and WildDeepfake (Wild) (Zi et al. 2020). See the supplementary material for more details.
@@
We employ CLIP-ViT-L (Radford et al. 2021) as the foundation model, which is pretraiend on LAION-400M. The input size of foundation model is $224\times 224$, we set $p=34$ for Input Transformations, so that trainable parameters of our method is 0.078M, which can be computed by Equation 2. We employed the AdamW optimizer with the learning rate 1.0, and the weight decay was fixed at 0. Besides, the data preprocessing transform is as same as the original CLIP(Radford et al. 2021) At last, the visual prompt $\boldsymbol{\delta}$ is initialized by zero padding.
@@
In this work, we mainly report the area under the ROC curve (AUC) to compare with prior works. The video-level results are obtained by averaging predictions over each frame on an evaluated video.

### v4

@@ Abstract
The proliferation of deepfake faces poses huge potential negative impacts on our daily lives. Despite substantial advancements in deepfake detection over these years, the generalizability of existing methods against forgeries from unseen datasets or created by emerging generative models remains constrained. In this paper, inspired by the zero-shot advantages of Vision-Language Models (VLMs), we propose a novel approach that repurposes a well-trained VLM for general deepfake detection. Motivated by the model reprogramming paradigm that manipulates the model prediction via input perturbations, our method can reprogram a pre-trained VLM model (e.g., CLIP) solely based on manipulating its input without tuning the inner parameters. First, learnable visual perturbations are used to refine feature extraction for deepfake detection. Then, we exploit information of face embedding to create sample-level adaptative text prompts, improving the performance. Extensive experiments on several popular benchmark datasets demonstrate that (1) the cross-dataset and cross-manipulation performances of deepfake detection can be significantly and consistently improved (e.g., over 88% AUC in cross-dataset setting from FF++ to WildDeepfake); (2) the superior performances are achieved with fewer trainable parameters, making it a promising approach for real-world applications.
@@ Introduction
Deepfake refers to a series of deep learning-based facial forgery techniques(Li et al. 2020a; Xu et al. 2022; Chen et al. 2024a) that can swap or reenact the face of one person in a video to another. In recent years, deepfake videos (a.k.a, deepfakes) have gained substantial attention due to their potential by creating and spreading false information. Thus, detecting deepfakes has emerged as a crucial research topic to reduce such security risks.
Existing methods treat deepfake detection as a binary classification problem and predominantly utilize CNNs (e.g., Xception or EfficientNet) as the backbones of classifier. In addition, some works (Qian et al. 2020; Li et al. 2020b; Zhao et al. 2021; Yan et al. 2023b, 2024b) propose to introduce auxiliary clues, including modalities (e.g., frequency) or supervision (e.g., forgery masks) information for learning subtle forgery artifacts. Despite these advancements, evaluating forgeries from unseen datasets and synthesized by unseen methods beyond the training data still poses a significant challenge for practical deepfake detection.
@@
Vision Language Models (VLMs), such as CLIP (Contrastive Language-Image Pre-training) (Radford et al. 2021), demonstrate robust zero-shot and few-shot generalization capabilities across diverse downstream tasks. To improve deepfake detector generalization, VLFFD (Sun et al. 2023) fully fine-tunes a CLIP-based model, suggesting the effectiveness of CLIP. However, excessive adjustment of parameters in pre-trained models can disrupt the great pre-trained knowledge, potentially leading to suboptimal performance. Thus, developing a cost-efficient approach is demanding.
Due to the inherent vulnerability of deep networks to adversarial attacks, a novel concept called model reprogramming (Elsayed, Goodfellow, and Sohl-Dickstein 2019; Chen 2024) has been proposed to re-purpose a well-trained model trained in a source domain to perform a target-domain task, just through learning universal input perturbations without modifying the source-domain model parameters. In other words, an additive offset applied to a deep network’s input would be sufficient to adapt the network to a new task without the need of re-training or fine-tuning its inner parameters. The reprogramming paradigm requires significantly fewer learnable parameters compared to fine-tuning numerous parameters within the model. Inspired by the above advancements, we advocate that reprogramming a well-trained foundation model to identify deepfakes is a promising way to improve generalization and training efficiency.
In this paper, we propose RepDFD, a novel method to reprogram a pre-trained CLIP model for effective and general deepfake detection. RepDFD solely learns task-specific visual perturbations (a.k.a, visual prompts) in pixel spaces for the deepfake detection task while keeping the entire CLIP model frozen. Specifically, we introduce Input Transformation to merge the image and the visual prompt and then feed it into the image encoder of CLIP. Furthermore, we propose Face2Text Prompts to generate text prompts merging information of face embedding, and then feed them into the text encoder of CLIP to guide the optimization of the visual prompts. This straightforward method enables the CLIP model to effectively detect deepfakes. Moreover, since the internal parameters are not trained, the foundation CLIP model can be reused for the other vision tasks. Extensive experiments on several popular deepfake benchmarks demonstrate that (1) the cross-dataset and cross-manipulation performances can be significantly and consistently enhanced by equipping RepDFD for a pre-trained CLIP model; and (2) the superior performances are achieved with fewer trainable parameters (see in Fig. 1), making it a promising approach for real-world applications.
@@
  • This is the first work to explore model reprogramming paradigm for deepfake detection tasks.
  • We have proposed RepDFD to reprogram a pre-trained CLIP model by solely processing its image and text inputs without tuning the inner parameters. Thus, RepDFD can be seamlessly adapted to other foundation models.
  • We have conducted extensive experiments on several benchmark datasets, and have demonstrated that RepDFD is general and efficiency for deepfake detection.
Related Work
@@
The past five years have witnessed a wide variety of methods proposed for defending against the malicious usage of deepfakes. Currently, the majority of deepfake detection methods are based on deep learning, leveraging generic CNNs (e.g., Xception, EfficientNet) as the backbones of the classifier. Furthermore, several works (Qian et al. 2020; Zhao et al. 2021) utilize frequency information or localize the forged regions to improve the performance of detectors. Nevertheless, the generalization challenge still hinders the application of deepfake detectors in real-world scenarios. To address such issue, several works (Li et al. 2020b; Shiohara and Yamasaki 2022; Nguyen et al. 2024; Lin et al. 2024b) introduce the augmented deepfake data, where two different faces are blended, during the training. However, all of the above methods typically involve retraining backbone networks. In general, employing more powerful backbone networks (e.g., replacing CNNs with ViTs or VLMs) produces better performances but at the cost of increased computational costs in the training stage.
@@
CLIP is a vision-language model (Radford et al. 2021), that is able to perform flexible zero-shot transfer to unseen classes using text prompts. Since CLIP is pre-trained to predict whether an image matches a textual description, it naturally fits zero-shot recognition. This is achieved by comparing image features with the classification weights synthesized by the text encoder, which takes as input textual descriptions specifying classes of interest. The CLIP model contains an image-encoder $E_{I}(\cdot)$ and a text-encoder $E_{T}(\cdot)$ such that the cosine similarity between the features $E_{I}(x_{k})$ and $E_{T}(t_{k})$ are maximized with respect to each pair $k$. Compared with the traditional classifier learning approach where closed-set visual concepts are learned from random vectors, vision-language pre-training allows open-set visual concepts to be explored through a high-capacity text encoder, leading to a broader semantic space and in turn making the learned representations more transferable to various vision tasks. Several works (Sha et al. 2023; Ojha, Li, and Lee 2023; Cozzolino et al. 2024) has explore that using CLIP model for detecting generated images by GANs and Diffusion Models. VLFFD (Visual-Linguistic Face Forgery Detection) (Sun et al. 2023) proposes to fully fine-tune CLIP with fine-grained text prompts for deepfake detection. In this work, we propose a cost-efficient method to adapt a pre-trained CLIP model for general deepfake detection.
@@
Figure 2: Overall framework of our proposed method. The core idea involves optimizing an universal visual prompt on a frozen CLIP model and generating sample-level text prompts (where the placeholder [FE] is replaced by a face embedding), aiming to adapt the model for the deepfake detection task.
Model Reprogramming
@@
Parameter Efficient Fine-Tuning (PEFT) methods (Tsao et al. 2024), which significantly reduce computational and storage overheads, have garnered increasing attention. By fine-tuning only a limited number of additional parameters, PEFT can adapt a large pre-trained model to achieve outstanding performance on targeted tasks. One of the PEFT methods, model reprogramming (Chen 2024), which incorporates prompts into inputs, offers an effective framework for repurposing models for various task-specific applications. The framework draws significant inspiration from adversarial reprogramming, which was first introduced by Elsayed et al (Elsayed, Goodfellow, and Sohl-Dickstein 2019). Model reprogramming aims to re-use and re-align the data representation, from an existing model, for a separate task without fundamental changes to the model’s inner parameters. This paradigm re-purposes existing knowledge by strategically transforming inputs and outputs, bypassing extensive inner model parameter fine-tuning. Reprogramming techniques have been widely applied in various tasks in the past few years (Wang et al. 2022; Cai et al. 2024). Visual prompting (VP) (Bahng et al. 2022) introduces a universal perturbation directly into the input data to facilitate task-specific fine-tuning while keeping the pre-trained model intact. Although these PEFT methods have developed in other research fields, the application of PEFT methods to face forgery detection remains largely unexplored. In this study, we introduce a novel approach to reprogramming a pre-trained CLIP model for deepfake detection using only a few number of learnable parameters, designed to leverage the great generalization abilities of the pre-trained Vision-Language Model (VLM).
@@
This paper proposes RepDFD, which reprograms the pre-trained vision language model CLIPs for deepfake detection without altering internal parameters. Given a well-trained CLIP model (including the image encoder $E_{I}$ and the text encoder $E_{T}$), it introduces visual and textual prompts on the inputs for adapting the frozen $E_{I}$ and $E_{T}$ to identify deepfakes. To achieve this, we propose two modules, Input Transformation and Face2Text Prompts, which process the image and text before inputting to $E_{I}$ and $E_{T}$, respectively. In this way, we harness the power of the model reprograming to guide the CLIP model to focus on the deepfake detection task while preserving its inner parameters and pre-trained knowledge. The overall framework is depicted in Fig. 2. In what follows, we elaborate on the details of Input Transformation, Face2Text Prompts and the optimization pipeline.
@@
RepDFD refines the visual features from the pre-trained CLIP model for deepfake detection tasks by introducing learnable visual prompts containing perturbations that improve the visual encoder’s features. A similar scheme like VP (Bahng et al. 2022) and AutoVP (Tsao et al. 2024) is adopted to initialize the visual prompt, placing it around images for input transformations. Input images are first resized smaller and subsequently incorporated with the visual prompts, aiming to match the input size of the pre-trained image encoder $E_{I}$. Fig. 3 demonstrates the details of Input Transformation. Given an original image $\mathbf{X}$, the input transformation aims to merge the universal visual prompt $\boldsymbol{\delta}$ to $\mathbf{X}$. Formally, the process of input transformations can be described as
$\mathcal{T}_{p}(\mathbf{X},\boldsymbol{\delta})=\text{Resize}_{p}(\mathbf{X})+\boldsymbol{\delta},$ (1)
where $p$ is the width of the visual prompt. Finally, the original image $X$ of size $H\times W$ is resized to $H^{\prime}\times W^{\prime}=(H-2p)\times(W-2p)$, ensuring the transformed image does not overlap with $\boldsymbol{\delta}$. The number of trainable parameters in this method is model-agnostic, which can be computed by
#$\#Para=3(HW-H^{\prime}W^{\prime})=6p(H+W)-12p^{2}.$ (2)
This expression indicates that $\boldsymbol{\delta}$ is applied to the RGB channels of the image. Therefore, for training or inference, $\mathcal{T}_{p}(\mathbf{X},\boldsymbol{\delta})$ is fed into image encoder $E_{I}$ to get the image feature $f$, i.e.,
$f=E_{I}(\mathcal{T}_{p}(\mathbf{X},\boldsymbol{\delta})).$ (3)
@@
To map model outputs to the target label, the CLIP model utilizes text prompts to align image features in a shared latent space, enabling classification by matching the closest text description to the image features. Therefore, a particular text prompt must be created for every class by injecting the information of the category. For instance, one of the official unified templates of text prompts is “A photo of a [cls]”, where [cls] is a placeholder of class label. However, in the context of deepfake detection, the binary class labels “real face” and “fake face” (or synonymous terms) lack specificity compared to labels such as “cat” or “dog”, making them difficult to accurately identify using existing text templates. To address this issue, VLFFD (Sun et al. 2023) introduces fine-grained text descriptions manually to fine-tune both $E_{I}(\cdot)$ and $E_{T}(\cdot)$ in the CLIP model. However, designing a hand-crafted text prompt that comprehensively describes forgery clues is difficult and labor-intensive.
In this work, we aim to design a simple but effective text prompts generation method to fully harness the generalizable capabilities of the CLIP model for deepfake detection. The success of personalized image generation methods (Chen et al. 2024b; Papantoniou et al. 2024) has shown that introducing pre-trained image embedding as textual embeddings is beneficial for maintaining the consistency of generated subjects. Considering that deepfakes involve the manipulation of faces, the inconsistency of facial representations between real and fake faces is a reliable clue. Inspired by this, we suggest exploiting the CLIP text encoder to integrate face embeddings into text embeddings. In this way, the sample-level face information, which is difficult to express by language, can be incorporated into textual prompts, thereby enhancing the CLIP model’s capability for efficient deepfake detection.
To achieve this, a Face2Text module is proposed and the details are shown in Fig. 4. The prompt template $\mathbf{T}$ is designed as “A [cls] photo of a [FE] person”. Given a face encoder $E_{F}$, we obtain a face embedding $S^{*}$ for an input image $\mathbf{X}$:
$\centering S^{*}=f_{\text{map}}(E_{F}(\mathbf{X})).\@add@centering$ (4)
where $f_{map}$ represents a frozen linear layer with a random initialized, projecting the face embeddings to align the input dimension of the text encoder $E_{T}$. By the operation $TTE$ (Text To Embedding), the text prompt $\mathbf{T}$ is initially tokenized and converted into word embeddings $\mathbf{t}$, which can be expressed as
$\centering\mathbf{t}=TTE(\mathbf{T}).\@add@centering$ (5)
Then, the placeholder token [FE] is replaced with the face embedding $S^{*}$, as shown below
$\centering\mathbf{t}=Face2Text(\mathbf{t},S^{*}).\@add@centering$ (6)
We also consider the plain prompt template “A [cls] photo of a person,” which is not processed by Eq. (6). Thus, we get four different prompt templates for real and fake labels (see Tab. 1). An interesting finding according to the results is that an asymmetric setting for real and fake categories ($\{T_{0},T_{3}\}$) achieves the best performance. The in-depth analysis is discussed later.
@@
Therefore, for training or inference, a text embedding $t$ is fed into text encoder $E_{T}$ of the pre-trained CLIP to get the text feature $w$, i.e.,
$w=E_{T}(\mathbf{t}),$ (7)
@@
$T_{0}$ “A real photo of a person”
$T_{1}$ “A fake photo of a person”
$T_{2}$ “A real photo of a [FE] person”
$T_{3}$ “A fake photo of a [FE] person”
@@
$\centering P(y_{i}\mid x)=\frac{\exp(\cos(w_{i},f)/\tau)}{\sum_{j=0}^{1}\exp(\cos(w_{j},f)/\tau)},i=\{0,1\}\@add@centering$ (8)
where $cos(\cdot)$ denotes the cosine similarity, $\tau$ is the temperature parameter of CLIP, $f$ is the image feature, and $w_{i}$ is the text feature from the text prompt of label $y_{i}$.
During training, the learnable visual prompt $\boldsymbol{\delta}$ is optimized by maximizing the likelihood of the correct label. During inference, $\boldsymbol{\delta}$ is pad around the shrunk test samples and for predictions. On the training dataset $\mathcal{D}$, the optimization target is to minimize $\mathcal{L}$ by tuning $\boldsymbol{\delta}$, which can be formulated as
$\centering\mathcal{L}=\mathbb{E}_{\mathbf{X}\sim\mathcal{D}}\left[-\sum_{n}^{N}P(y_{n}|x_{n})\log(P(y_{n}|x_{n}))\right],\@add@centering$ (9)
where $N$ is the size of training dataset $\mathcal{D}$, $x_{n}$ is the $n$-th image sample, and $y_{n}$ is the ground truth label for $x_{n}$. We initialize the $\boldsymbol{\delta}$ to zero, and follow a simple gradient-based approach to directly optimize the visual prompt via back-propagation which updates $\boldsymbol{\delta}$ at step $k$ by
$\centering\boldsymbol{\delta}^{k+1}=\boldsymbol{\delta}^{k}-\gamma{\nabla_{\boldsymbol{\delta}^{k}}\mathcal{L}},\@add@centering$ (10)
@@
Notably, Eq. (8) has a similar formulation to a traditional fully connected layer with input dimension $N$ and output dimension $2$ if we treat the text features $[w_{0},w_{1}]$ as the ‘weight’ of a fully connected layer for the real and fake categories. Given that the adaptative text feature for each input image in our method, it implies that the model will classify samples based on sample-level dynamic classifiers. Therefore, the Face2Text module will also significantly enrich the supervised information during optimization.
@@
SLADD (CVPR 2022) 79.70 - - 77.20 SBI (CVPR 2022) 88.61 70.27 84.80 71.70
@@
Following most previous works, we mainly conducted training on the FaceForensics++ (FF++) (Rossler et al. 2019). It contains 1000 Pristine (PT) videos (i.e., the real sample) and 5000 fake videos forged by five manipulation methods, i.e., Deepfakes (DF), Face2Face (F2F), FaceSwap (FS), NeuralTextures (NT) and FaceShifter (FSh). Besides, FF++ provides three quality levels in compression for these videos: raw, high-quality (HQ) and low-quality (LQ). The HQ version of FF++ is adopted by default in this paper. The samples were split into disjoint training, validation, and testing sets at the video level follows the official protocol. To demonstrate the performances in cross-dataset settings, four additional datasets are adopted, i.e., Celeb-DF-v2 (CDF) (Li et al. 2020c), DeepFake Detection Challenge preview (DFDCP) (Dolhansky et al. 2019), DeepFake Detection Challenge public (DFDC) (Dolhansky et al. 2020) and WildDeepfake (Wild) (Zi et al. 2020).
@@
We employed CLIP-ViT-L (Radford et al. 2021) as the foundation model, which is pretraiend on LAION-400M. A pre-trained Transface (Dan et al. 2023) was employed as $E_{F}$. The size of input for foundation model is $224\times 224$. We set $p=34$ for Input Transformations, so that trainable parameters of our method is 0.078M according to Eq. (2). We employed the AdamW optimizer with the learning rate 1.0, and the weight decay was fixed at 0. Besides, the data preprocessing transform was as same as the original CLIP (Radford et al. 2021) Notably, the visual prompt $\boldsymbol{\delta}$ was initialized by zero.
@@
In this work, we mainly report the area under the ROC curve (AUC) to compare with prior works. The video-level results are obtained by averaging predictions over each frame on an evaluated video. To facilitate a comprehensive comparison of our method with others, we also present the results of the equal error rate (EER) in our appendix (Lin et al. 2024a).